OpenStudy (anonymous):

What is the missing term, x, in the pattern below? 1, 16, x, 256, 625, … A. 81 B. 125 C. 128 D. 136

OpenStudy (anonymous):

What is the pattern? Seems like x^n but we don't know n. Explore fourth power of 1,2,3,4,5 16 = 2^4 256 = 4^4 625 = 5^4 so x is 3^4 = ?

OpenStudy (anonymous):

81?

OpenStudy (anonymous):

Yes i think it is 81.